Tom, Ted, Tony, and Terry worked together on a job. The Job paid$500. Each boy was paid according to the amount of work he contributed to the job. Ted and Tony earned $280 together. Ted and Tom earned $260 together. Ted and Terry got $220 together. What percent of the job did each boy do?

THIS PROBLEM IS NOT FROM A TEXTBOOK

Answer by stanbon(48545)

Tom, Ted, Tony, and Terry worked together on a job. The Job paid$500. Each boy was paid according to the amount of work he contributed to the job. Ted and Tony earned $280 together. Ted and Tom earned $260 together. Ted and Terry got $220 together. What percent of the job did each boy do?
===============
Let d be Ted's portion; y be Tony's ; r be Terry's ; m be Tom''s
d+y+r+m=500
d+y+0+0=280
d+0+0+m=260
d+0+r+0=220
-----------
Using the Matrix function of a TI calculator to solve this system
of equations I get:
Ted = $130 ; so Ted did 130/500 = 26%
Tony= $150 ; so Tony did 150/500= 30%
Terry=$90 ; so Terry did 90/500= 18%
Tom = $130 ; so Tom did 130/500 = 26%
==================
Cheers,
Stan H.